# AI for Accelerated Materials Design

### v3

@@ Appendix B Batch and pad algorithms
$B_{n}\leftarrow\text{next-multiple-64}(n*B_{G}/N)$ {padding target}
$B_{e}\leftarrow\text{next-multiple-64}(e*B_{G}/N)$
return $B_{n}$, $B_{e}$, $B_{G}$
Appendix C Timing experiments setup
@@
Appendix D Mean versus median in timing measurements
@@
Appendix E CPU only timing results
The algorithm performances are different when running only on CPU. The batching (inclusive of padding) performance is the same as when running on a system that has both a GPU and a CPU which indicates that the batching is executed only on CPU. We experimented with trying to run part of the batching and padding code on GPU using JAX commands but found no speedup. The gradient-update step, however, is much slower on CPU. The mean timing results, from ten iterations of one hundred thousand training steps, are shown in Fig. 10 for the MPEU model. The median is shown in Fig. 11 for the same model. We ran one hundred thousand training steps instead of two million training steps since the computer cluster we used had a time limit of twelve hours for experiments. From the mean results, we see the dynamic batching algorithm is fastest. The median results, however, show that when the effect of recompilations are reduced, the static-64 algorithm is the fastest except for batch size 128 on the AFLOW dataset. This suggests that for longer training times the static-64 algorithm will be the fastest. The results for the SchNet model can be seen in the accompanying code repository.
@@ Appendix F Multi-GPU behavior
We used JAX parallel map (pmap) function to run the batching algorithms on a multi-GPUs of four GPUs. In this case, we saw that the static-64 and static-$2^{n}$ algorithms did not work out of the box with pmap since the algorithm expects the batches sent to each GPU to be the same size. For the AFLOW dataset, we saw roughly that the four batches had different shapes after padding when using the static-$2^{n}$ algorithm roughly 90% of the time, and close to 99% when using the static-64 algorithm. It’s possible to alter the static-64 and static-$2^{N}$} algorithm to pad each batch of data (e.g. for four GPUs there are four batches) using the largest number of nodes and edges across all batches. The results of such tests, however, will be highly implementation dependent, more-so than in the single-GPU case. To simplify the comparison of the algorithms, we evaluate the algorithms only on single-GPU setups and note that the dynamic algorithm is easily ported to a multi-GPU setup using JAX’s pmap function.
@@ Appendix G Test metric results
The test metric (RMSE) curves for each of the ten iterations are averaged for each batch size, model and dataset combination. The resulting mean curve is shown for the MPEU and SchNet models using the QM9 data in Fig. 12. We do not see significant differences in the test performance for the static-$2^{N}$ or dynamic batching algorithm. Note that we do not expect nor do we see any noticeable difference in learning from the static-64 to static-$2^{N}$ algorithms since the difference between the methods is the padding scheme which does not affect the loss. Therefore, the static-64 curve is left out of Fig. 12 for visual clarity. The results for AFLOW can be seen in the accompanying code repository.
Figure 12: Mean test RMSE curves for the MPEU model (left) and SchNet (right) on QM9 test for batch sizes of 16, 32, 64 and 128 (from top to bottom).
@@ Appendix H Static-constant batching
static-2N AFLOW SchNet 32 0.71 1.38 2.01
@@ Appendix I Student’s t-tests on test metrics
The Student’s t-test was performed in a pairwise fashion between two batching algorithms upon their distribution of final test RMSE values. For the SchNet model on the QM9 dataset for batch size 64, a heatmap of the t-test values is shown in Fig. 13 and the associated p-values in Fig. 14.
@@
Figures for other datasets and models can be found in the Software and Data section.

### v4

@@ Appendix A Batch and static batching algorithms
return $\textsc{Graph}(\mathbf{X}_{all},\mathbf{E}_{all},\mathbf{u}_{all},\mathbf{s}_{all},\mathbf{r}_{all},\mathbf{n}_{v},\mathbf{n}_{e})$
The static-constant batching algorithm first finds the graph with the largest number of nodes and the graph with the largest number of edges. These values are saved and when rounded to the nearest multiple of 64 serve as targets for padding. The static batching algorithm is shown in Algorithm 5. The algorithm grabs a set of batch size minus one graphs, and then pads the set of graphs to the next multiple of 64 (static-64) or next power of two (static-$2^{N}$). Note that Algorithm 5 shows the static-$2^{N}$ implementation.
Algorithm 5 Static batching
Input: Dataset $\mathcal{G}$, Start index $s$, Batch size NN
$\mathcal{B}\leftarrow\emptyset$ {Initialize batch container}
$\Sigma_{v}\leftarrow 0,\quad\Sigma_{e}\leftarrow 0$
 {Collect N1N-1 real graphs (leaving 1 slot for padding)}
for $k=0$ to N2N-2 do
  $G_{i}\leftarrow\mathcal{G}[s+k]$
  $\mathcal{B}\leftarrow\mathcal{B}\cup\{G_{i}\}$
  $\Sigma_{v}\leftarrow\Sigma_{v}+|V_{i}|$
  $\Sigma_{e}\leftarrow\Sigma_{e}+|E_{i}|$
end for
 {Pad to nearest power of 2 (or multiple of 64)}
$T_{v}\leftarrow\textsc{NextPower2}(\Sigma_{v})$
$T_{e}\leftarrow\textsc{NextPower2}(\Sigma_{e})$
$T_{g}\leftarrow N$
$T\leftarrow(T_{v},T_{e},T_{g})$
return Batch(PadToTarget($\mathcal{B},T$))
Appendix B Timing experiments setup
@@
Appendix C Mean versus median in timing measurements
@@
Appendix D Profiling behavior as a function of the number of training steps run
The running average combined time per step is shown in Fig. 11 on the AFLOW dataset for the MPEU model using a batch size of 32. The dynamic algorithm is initially the fastest. The static-64 algorithm is initially significantly slower due to the large number of recompilations that happen in the first 100,000 training steps. After 300,000 training steps, the effect of the recompilations on the static-64 algorithm’s performance is reduced and the static-64 and dynamic methods are roughly equally as fast.
Figure 11: The running average of the combined time (sum of the batching step and gradient-update step) required per training step as a function of the total number of training steps run. Here only a single iteration is run for the batch size 32, MPEU model and AFLOW dataset for the dynamic, static-64 and static-$2^{N}$ algorithms.
Appendix E CPU-only timing results
This section provides supplementary details regarding the CPU only timing experiments. The algorithm performances are different when running only on CPU. The batching (inclusive of padding) performance is the same as when running on a system that has both a GPU and a CPU, which indicates that the batching is executed only on CPU. We experimented with trying to run part of the batching and padding code on GPU using JAX commands but found no speedup. The gradient-update step, however, is much slower on CPU. The mean timing results, from ten iterations of one hundred thousand training steps, are shown in Fig. 12 for the MPEU model. The median is shown in Fig. 13 for the same model. We ran one hundred thousand training steps instead of two million training steps since the computer cluster we used had a time limit of twelve hours for experiments. From the mean results, we see the dynamic batching algorithm is fastest. The median results, however, show that when the effect of recompilations are reduced, the static-64 algorithm is the fastest except for batch size 128 on the AFLOW dataset. This suggests that for longer training times the static-64 algorithm will be the fastest. The results for the SchNet model can be seen in the accompanying code repository.
@@ Appendix F Multi-GPU behavior
We utilized JAX’s parallel map (pmap) function to assess the feasibility of the algorithms on a multi-GPU setup comprising four A100 GPUs. The dynamic batching algorithm is natively compatible with distributed setups; JAX’s pmap distributes the graph batches across devices without requiring changes to the batching logic itself. The static-64 and static-$2^{n}$ algorithms, however, did not work out of the box with pmap since the algorithm expects the batches sent to each GPU to be the same size. For the AFLOW dataset, we saw that the four batches for each GPU had different shapes after padding when using the static-$2^{n}$ algorithm roughly 90% of the time, and close to 99% when using the static-64 algorithm. It’s possible to alter the static-64 and static-$2^{N}$} algorithm to pad each batch of data using the largest number of nodes and edges across all batches. The results of such tests, however, will be highly implementation dependent. To simplify the comparison we evaluate the algorithms on single-GPU setups.
@@ Appendix G Test metric results
The mean test metric (RMSE) values as a function of the number of training steps are averaged across each of the ten iterations are averaged for each batch size, model and dataset combination and are shown in Fig. 14 for the MPEU and SchNet models using QM9 data. We do not see significant differences in the test performance for the static-$2^{N}$ or dynamic batching algorithm. The results for the AFLOW test dataset can be seen in Fig. 15. Note that we do not expect nor do we see any noticeable difference in learning between the static-64 and static-$2^{N}$ algorithms since the difference between the methods is the padding scheme, which does not affect the loss. Therefore, the static-64 RMSE curve is left out of the two figures.
Figure 14: Mean RMSE values as a function of the number of training steps for the MPEU model (left) and SchNet (right) on QM9 test data for batch sizes of 16, 32, 64 and 128 (from top to bottom).
Figure 15: Mean RMSE values as a function of the number of training steps for the MPEU model (left) and SchNet (right) on AFLOW test data for batch sizes of 16, 32, 64 and 128 (from top to bottom).
@@ Appendix H Static-constant batching
static-$2^{N}$ AFLOW SchNet 32 0.71 1.38 2.01
@@ Appendix I Student’s t-tests on test metrics
For completeness, we also provide pairwise Student’s t-test metrics on the final RMSE test loss. Note, the Student’s t-test works best when the underlying distributions are normal Ross (2017), which is a strong assumption. For this reason, in the main text the Mann–Whitney U test is used and this approach is included for completeness. One can either use the t-test statistic (critical value approach) or the associated p-value to determine whether to reject the null hypothesis Hartmann et al. (2023) that the test metrics after two million training steps from different batching methods belong to the same underlying distribution. We adopt p-value criteria from the literature Ross (2017), that argues that there is little evidence that the distribution of the test metrics differ significantly if the p-values are greater than 0.10. For p-values between 0.05 to 0.10, there is moderate evidence that the distributions are different. For lower p-values the evidence is strong. Ignoring the p-values that are greater than 0.10, we can look at several interesting results. For the QM9 data, SchNet model, and batch size 64, the static-$2^{N}$ and dynamic methods have a p-value of 0.033. For the QM9 data, SchNet model, and batch size 32, the static-$2^{N}$ and static-64 algorithms have p-values with the dynamic algorithm of 0.067 and 0.053, respectively. For QM9 data, MPEU model, and batch size 64, the static-$2^{N}$ and dynamic methods have a p-value of 0.071. These results show that for some datasets, models, and batch sizes, the dynamic algorithm gives significantly different test metrics than the static algorithm. For the majority of combinations, however, the results are similar, which may be due to use of the Adam optimizer whose adaptive estimation of first and second moments may reduce the effect of occasionally having smaller batch sizes. Note, that unlike in the main text, a Bonferroni correction is not used in this analysis, which means we are not scaling the p-value criteria based on the number of pairwise tests that are run.
@@
Fig. 16 shows a heatmap of pairwise t-test values and Fig. 17 shows the associated p-values across batching methods for the SchNet model, batch size 64, and QM9 dataset . Fig. 18 and Fig. 19 show a pairwise t-test heatmap of values and associated p-values respectively for the SchNet model, batch size 64 and AFLOW dataset. The complete set of plots for the other combinations of datasets, models and batch sizes can be found in the Data Availability section.
Figure 18: Heatmap of pairwise Student’s t-test values on the distribution of test RMSE values for the SchNet model, AFLOW data and a batch size of 64.
Figure 19: Heatmap of p-values from the pairwise Student’s t-test on the distribution of test RMSE values for the SchNet model, AFLOW data and a batch size of 64.
Appendix J Mann-Whitney U-Statistic.
We show here a pairwise heatmap for the statistic values in Fig. 20 and associated p-values in Fig 21. The complete set of plots for the other combinations of datasets, models and batch sizes can be found in the Data Availability section.
Figure 20: Heatmap of pairwise Mann-Whitney U test values on the distribution of test RMSE values for the SchNet model, AFLOW data and a batch size of 16.
Figure 21: Heatmap of pairwise Mann-Whitney U test associated p-values with a Bonferroni correction on the distribution of test RMSE values for the SchNet model, AFLOW data and a batch size of 16.